OpenStudy (erich):

y=x-1 x-y^2=9 Solve for X

OpenStudy (erich):

This is how Ive solved it so far... x-(x-1)^2=9 x-x^2-2x+1=9

OpenStudy (erich):

-x^2-x=8

OpenStudy (erich):

x^2=x-8

OpenStudy (erich):

so x= the sqrt of x-8?

OpenStudy (anonymous):

x-(x-1)^2=9 x-(x^2-2x+1)=9 x-x^2+2x-1 = 9 x^2 -3x + 10 = 0 \[x = 1.5\pm2.78i\]

OpenStudy (erich):

i dont understand how that is the answer